Grade 12Mechanics

A600kg rocket is set for a vertical firing, if the exhaust speed is1000m/S, the mass of the gas ejected per second to supply the thrust needed to overcome the weight of rocket is

Sri
7 Years agoGrade 12

1 Answer

Khimraj
7 Years ago
F = mg = v*dm/dt
dm/dt = mg/v
= 600*9.8/1000 = 5.88 kg/s
Hope it clears..........................................